Suppose that Gyp Sum Industries currently has the balance sheet shown below, and that sales for the year just ended were $10.9 million. The firm also has a profit margin of 25 percent, a retention ratio of 30 percent, and expects sales of $8.9 million next year.

Assets Liabilities and Equity
Current assets $ 2,621,000 Current liabilities $ 2,557,140
Fixed assets 4,900,000 Long-term debt 1,950,000
Equity 3,013,860
Total assets $ 7,521,000 Total liabilities and equity $ 7,521,000

If all assets and current liabilities are expected to shrink with sales, what amount of additional funds will Gyp Sum need from external sources to fund the expected growth? (Enter your answer in dollars not in millions. Negative amount should be indicated by a minus sign.)
Answer #1

Total Assets = $ 7521000, Current Liabilities = $ 2557140, Current Sales = $ 10.9 million and Expected Sale = $ 8.9 million

% Decrease in Sales = (8.9 - 10.9) / 10.9 ~ - 0.1835 or -18.35 %

As current liabilities and total assets are expected to shrink proportionally with sales, new total assets = (1-0.1835) x 7521000 ~ $ 6141000 and Current Laibilities = (1-0.1835) x 2557140 ~ $ 2087940

Profit Margin = 25 % of Sales

Net Income = 0.25 x 6141000 = $ 1535250 and Retention Ratio = 0.3

Retained Earning = Retention Ratio x Net Income = 0.3 x 1535250 = $ 460575

Additional Funds Needed = New Total Assets - New Current Liabilities - Long-Term Debt - Retained Earnings - Equity = 6141000 - 2087940 - 1950000 - 460575 - 3013860 = $ -1371375